Esperanto Corporation wants to determine a linear function relating price, p(x), to demand where x = number of doodads sold. They sold 900 doodads per week when price was 94 dollars for each doodad. The number of doodads sold dropped to 400 doodads per week when price was raised to 104 dollars for each doodad. Write the function that models price as a function of number of doodads sold

Answer by stanbon(75887)

Esperanto Corporation wants to determine a linear function relating price, p(x), to demand where x = number of doodads sold.
They sold 900 doodads per week when price was 94 dollars for each doodad.
The number of doodads sold dropped to 400 doodads per week when price was raised to 104 dollars for each doodad.
----------------------------------------------
Write the function that models price as a function of number of doodads sold
---
You have two points: (900,94) and (400,104)
slope = (104-94)/(-500)= -1/50
---
The linear form is y = mx + b
104 = (-1/50)*400 + b
b = 104 + 8
b = 112
----------------
Linear Equation:
p(x) = (-1/50)x + 112
===============================
Cheers,
Stan H.